### v2

# Exact $n+2$ Comparison Complexity for the N-Repeated Element Problem

Abstract
This paper establishes the exact comparison complexity of finding an element repeated $n$ times in a $2n$-element array containing $n+1$ distinct values, under the equality-comparison model with $O(1)$ extra space.
We present a simple deterministic algorithm performing exactly $n+2$ comparisons and prove this bound tight: any correct algorithm requires at least $n+2$ comparisons in the worst case.
The lower bound follows from an adversary argument using graph-theoretic structure. Equality queries build an inequality graph $I$; its complement $P$ (potential-equalities) must contain either two disjoint $n$-cliques or one $(n+1)$-clique to maintain ambiguity. We show these structures persist up through $n+1$ comparisons via a “pillar matching” construction and edge-flip reconfiguration, but fail at $n+2$.
This result provides a concrete, self-contained demonstration of exact lower-bound techniques, bridging toy problems with nontrivial combinatorial reasoning.
Keywords: comparison complexity; N-repeated element problem; equality comparison model; adversary argument; inequality graph; lower bound.
1 Introduction
This paper introduces lower-bound techniques through a concrete, easily stated example. The motivating problem is inspired by LeetCode 961 [1].
Consider an array of length $2n$, containing exactly $n+1$ distinct values. Among these, one element appears $n$ times, while every other element appears once.
The task is algorithmically simple—it serves only as a vehicle to explore how optimality can be proved.
Several natural approaches illustrate the trade-offs between time and space:
  • A hash-table approach detects collisions in expected linear time but requires linear extra space.
  • Sorting the array and examining its middle segment achieves $O(1)$ extra space but incurs $O(n\log n)$ time.
  • Randomized pairwise comparisons often succeed quickly but may, in the worst case, fail to terminate.
In this work, I focus on a simple deterministic algorithm and prove its exact optimality.
The algorithm proceeds as follows. Partition the array into $n$ disjoint pairs and compare each pair for equality. If any equality is found, return that value immediately. Otherwise, each pair must contain exactly one copy of the majority element. The algorithm then performs two additional “local” comparisons: it tests whether one element of the first pair equals either element of the second pair; if neither test succeeds, the algorithm returns the remaining candidate.
This method performs exactly $n+2$ comparisons and uses only $O(1)$ additional space.
The natural question is whether this bound is optimal.
Under the equality-comparison model with constant auxiliary space, I prove that any correct algorithm must perform at least $n+2$ comparisons. Consequently, the above procedure achieves the exact comparison complexity—tight with no additional gap.
This result provides a concise yet rigorous demonstration of how lower bounds emerge from combinatorial structure, even in seemingly simple computational settings.
2 Computational Model
To justify a statement such as “Any correct algorithm must perform at least $n+2$ comparisons,” we must first define precisely what counts as an algorithm. Otherwise, one could imagine a nonsensical “oracle algorithm”:
I hand the array to a deity, and the deity immediately returns the correct answer.
The purpose of a computational model is to exclude such oracles while retaining all legitimate algorithms. In general, a good model should be:
  • broad enough to capture all reasonable algorithms; and
  • restrictive enough to make universal lower-bound proofs possible.
Although a Turing machine can faithfully represent any algorithm, it is too general for the kind of exact complexity reasoning pursued here. Instead, this paper adopts a more specific model: algorithms operate solely through equality comparisons.
The allowed operations are as follows:
  • The algorithm may select any pair of array elements to compare for equality.
  • It learns whether the two values are equal or different.
  • It may record this information, either explicitly in a small data structure or implicitly as part of its control logic (as in the optimal algorithm from the introduction).
  • It may adapt future comparisons based on previous outcomes and eventually output the repeated element.
In the original LeetCode formulation, the array elements are numbers, which in principle allows many additional operations: comparisons by magnitude, bitwise manipulations, or arithmetic tests. None of these are permitted in this model.
This restriction is intentional—it simplifies the logical structure of the lower-bound proof. Allowing richer operations would make the model more expressive but also render such proofs substantially more difficult, as will be discussed later.
3 Adversary Argument
The previous section defined the space of allowable algorithms. We now turn to the main question: how can one assert that any correct algorithm must perform at least $n+2$ comparisons?
Equivalently, the task is to show that every algorithm making at most $n+1$ comparisons fails on at least one valid input instance.
This form of argument is typically established through an adversary method [2]. We imagine a two-player game between the algorithm and an adversary, similar to a chess match. At each step, the algorithm chooses two elements to compare, and the adversary must respond immediately with either “equal” or “not equal.” Eventually, the algorithm must commit to an answer—its guess for the repeated element. The adversary wins if it can produce a concrete input consistent with all previous answers yet for which the algorithm’s output is incorrect.
If such an adversary can always force a win whenever the algorithm performs at most $n+1$ comparisons, then the lower bound is proved.
Thus, our goal is to construct a universal adversary that guarantees a winning response under this constraint.
The key idea is that the adversary maintains ambiguity by keeping multiple consistent input configurations possible. If, after answering all comparison queries, the adversary can still exhibit at least two distinct valid instances—each consistent with the observed answers but having different majority elements—then the algorithm necessarily fails on one of them. This property forms the foundation of the lower-bound proof.
4 Strategy: Graph Interpretation
How can the adversary maintain multiple consistent answers? Initially, before any comparisons have been made, all $2n$ array positions are unconstrained: any subset of $n$ elements could correspond to the majority element.
As the algorithm performs equality queries, the adversary naturally responds “not equal.” Each such answer imposes a new restriction—the two queried elements cannot both belong to the same majority group.
To reason about these restrictions systematically, we define two complementary graphs.
Let $I$ denote the inequality graph. It has a vertex for each array position and an edge between two vertices if and only if the algorithm has tested them for equality. Because the adversary always replies “not equal,” each comparison adds an edge to $I$.
Let $P$ denote the potential-equality graph, defined as the complement of $I$. An edge in $P$ indicates that the corresponding pair of elements could still be equal—that is, their equality has not yet been ruled out.
The majority element occurs exactly $n$ times, and all occurrences of that element are mutually equal. Hence, the set of positions containing the majority element forms an $n$-clique within $P$.
For the adversary to preserve ambiguity, $P$ must contain more than one possible $n$-clique. However, not all such cliques guarantee a win. If every $n$-clique in $P$ shares at least one common vertex, the algorithm can safely choose that vertex’s element as its output, since it is consistent with all configurations.
Therefore, the adversary wins precisely when the $n$-cliques of $P$ do not all intersect. In that case, at least two disjoint possibilities remain for the location of the majority element, ensuring that any algorithm making the current set of comparisons cannot yet be correct.
There are two simple structural conditions that guarantee non-intersection:
  1. 1.
    Two disjoint $n$-cliques.
    If $P$ contains two vertex-disjoint $n$-cliques, the adversary can pick either as the majority group, contradicting any fixed output of the algorithm.
  2. 2.
    One $(n+1)$-clique.
    An $(n+1)$-clique contains $n+1$ distinct $n$-cliques as its subsets, and no single vertex belongs to all of them. Thus, even in this configuration, there is no uniquely guaranteed majority candidate.
These two cases capture exactly the structural scenarios that enable the adversary’s success, and they form the foundation for the tight $n+2$ comparison lower bound established later.
5 Finding a Large Clique
To apply the previous reasoning, we must ensure that the required $n$- or $(n+1)$-cliques actually exist in the potential-equality graph $P$. How can such cliques be guaranteed?
The key insight comes from viewing $P$ in terms of its complement $I$, the inequality graph. Let $C$ denote a minimum vertex cover [3] of $I$. If the cover $C$ is small—that is, if |C|<n|C|<n—then the vertices outside the cover, $V(I)\setminus C$, induce a complete subgraph in $P$. Indeed, no edge of $I$ connects any two of these vertices, so all possible edges among them exist in $P$. Hence $P$ contains a clique of size
$|V(I)\setminus C|=2n-|C|>n,$
which suffices to guarantee the adversary’s ambiguity.
To bound the size of $C$, we rely on a simple combinatorial observation.
Lemma 5.1.
Let $G$ be a graph with $k$ edges. Then its minimum vertex cover satisfies $|C|\leq k$.
Proof.
Select one endpoint from each edge of $G$; this set of at most $k$ vertices meets every edge, and thus forms a valid vertex cover. Since a minimum cover cannot be larger, it follows that $|C|\leq k$. ∎
Applying this lemma to $I$, if the algorithm performs fewer than $n$ equality comparisons, then $I$ has fewer than $n$ edges. Consequently, its minimum vertex cover has size $|C|<n$, and as shown above, $P$ contains a clique of size at least $n+1$. Such a structure ensures the existence of multiple consistent instances, implying that any algorithm performing fewer than $n$ comparisons cannot be correct.
6 Finding Disjoint Cliques
What happens when the algorithm performs exactly $n$ comparisons? The vertex-cover lemma now yields $|C|\leq n$, so we can no longer guarantee an $(n+1)$-clique in $P$. Instead, we seek the second winning condition: two disjoint $n$-cliques in $P$.
We first strengthen the vertex-cover observation with a structural characterization.
Lemma 6.1.
Let $G$ have a minimum vertex cover of size exactly $k$. Then the edges of $G$ form a perfect matching [4]: no two edges share a common vertex.
Proof.
Suppose two edges share a vertex $v$. Then the set consisting of $v$ plus one endpoint from each remaining k2k-2 edges forms a vertex cover of size at most k1k-1, contradicting minimality. Thus, the edges must be vertex-disjoint. ∎
This structure admits a vivid geometric interpretation. Visualize the $n$ edges of $I$ as $n$ “pillars”: each pillar connects one vertex on the “table” (bottom level) to one vertex “above” (top level). Since the edges form a perfect matching, there are exactly $n$ bottom vertices and $n$ top vertices, with no sharing.
In the complement graph $P$, the absence of edges within each level implies complete connectivity within each level. Thus:
  • The $n$ bottom vertices form an $n$-clique in $P$.
  • The $n$ top vertices form another $n$-clique in $P$.
Moreover, these two cliques are vertex-disjoint by construction.
The adversary can now choose either clique as the majority set—perfectly consistent with all comparison answers but forcing any fixed algorithm output to fail on at least one possibility.
Table levelAbove levelII pillarsBottom $n$-clique ($P$)Top $n$-clique ($P$)
Figure 1: Two disjoint $n$-cliques in $P$ from a perfect matching in $I$. Blue pillars represent “not equal” edges; green edges represent cliques in $P$.
7 The $(n+1)$-Comparison Case
What happens when the algorithm performs exactly $n+1$ comparisons? After the first $n$ comparisons, we have the pillar structure with two disjoint $n$-cliques in $P$ (see Figure 1). Can the adversary preserve ambiguity?
Consider the (n+1)(n+1)st comparison. There are two cases:
  1. 1.
    Inter-layer comparison.
    If the algorithm compares across layers (bottom to top), the adversary responds “not equal.” All $n+1$ edges in $I$ are now inter-layer, so both layers remain complete $n$-cliques in $P$.
  2. 2.
    Intra-layer comparison.
    Suppose the algorithm compares two bottom vertices $b_{i}$ and $b_{j}$. The adversary responds “not equal” (adding $\{b_{i},b_{j}\}$ to II).
    The adversary then flips the pillar above $b_{i}$: swap so $b_{i}$ connects to $t_{j}$ and $b_{j}$ connects to $t_{i}$. Now the edge $\{b_{i},b_{j}\}$ becomes an inter-layer edge in this new configuration.
In both cases, the adversary ensures all $n+1$ edges of $I$ connect different layers. Thus $P$ retains two vertex-disjoint $n$-cliques: the bottom layer and the top layer. Ambiguity persists.
Only on the (n+2)(n+2)nd comparison does the adversary lose—it can no longer maintain this inter-layer property across both layers simultaneously.
8 Discussion
The key property enabling this exact $n+2$ bound is the restriction to equality comparisons only. This model captures only direct observations—each comparison adds precisely one edge to the inequality graph $I$, with no additional inferences possible.
What happens if we allow magnitude comparisons (<,=><,=\,>)? The transitivity of the total order fundamentally changes the proof structure. Suppose the algorithm compares $a_{i}<a_{j}$ and $a_{j}<a_{k}$; it can immediately infer $a_{i}<a_{k}$ without an explicit comparison. These transitive closures effectively add “free” edges to $I$, dramatically accelerating clique destruction in $P$.
For example, a single chain of $n$ magnitude comparisons could establish a total order over $n+1$ elements, forcing all but one to be distinct from the majority—potentially solving the problem in far fewer than $n+2$ direct comparisons.
This explains why equality-only algorithms require exactly $n+2$ comparisons, while numeric algorithms can do better by leveraging order transitivity. The equality model thus serves as an ideal “toy setting” for teaching lower-bound techniques: simple enough to analyze exhaustively, yet rich enough to demonstrate nontrivial combinatorial structure.
Acknowledgements
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.